# On the Hyperparameter Loss Landscapes of Machine Learning Algorithms

### v1

@@ 3 Experimental Setup
Table 2: Summarization of meta-information of our empirical study.
  Models Datasets Fidelities Summarization
 Model
Total
HPs
Total
Configs.
Cat.
Class.
Cat.
Reg.
Num.
Class.
Num.
Reg.
Image
Class.
Training
Data
Training
Epochs
Total
Configs.
# Land-
-scapes
XGB $5$ $14,960$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $2.56$M $342$
RF $6$ $11,250$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $1.92$M $342$
LGBM $5$ $13,440$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $2.30$M $342$
DT $5$ $24,200$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $4.14$M $342$
CNN $8$ $6,480$ $6$$\{0.1,0.25,\text{all}\}$$\{10,25,50\}$ $0.35$M $108$
@@
  Total (Before accounting $5$-fold cross-validation): $\bm{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}11.15\textbf{M}}}$ $\bm{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1,500}}$
Table 2 summarizes the meta-information of our empirical study, while detailed HP search space of each model and the principles we follow in designing them, are left in Appendix F.1. We first consider decision tree (DT(Safavian & Landgrebe, 1991) and three types of its ensembles: random forest (RF(Breiman, 2001), XGBoost (Chen & Guestrin, 2016) and LightGBM (Ke et al., 2017). We analyze the HP space of these models using the tabular benchmark proposed in Grinsztajn et al. (2022), which comprises $25$ regression and $32$ classification tasks (see Appendix F.2). These datasets span a broad range of complexities in terms of number of instances and features and are thus idea choice for comprehensive inspection of landscape characteristics. In addition to these, we also study convolutional neural networks (CNNs(Krizhevsky et al., 2012) on six classic image classification tasks (Appendix F.2) using a joint architecture and hyperparameter search (JAHS) (Bansal et al., 2022) space. We additionally consider another JAHS scenario, for which we adopt the NASBench-HPO (Klein & Hutter, 2019) data included in HPOBench (Eggensperger et al., 2021). This includes $62,208$ configurations of a feed-forward neural network (FCNet) evaluated on $4$ UCI datasets
For each dataset, unless predefined, we randomly split the data into training ($80\%$) and test ($20\%$) set. For all HP configurations $\bm{\lambda}\in\bm{\Lambda}$ of each model, we exhaustively evaluate $\mathcal{L}(\bm{\lambda})_{\text{train}}$ and $\mathcal{L}(\bm{\lambda})_{\text{test}}$ using $5$-fold cross-validation. Here, we use root mean squared error (RMSE) and $R^{2}$ score to serve as the loss function $\mathcal{L}$ for regression tasks, and for classification, we use accuracy and ROC-AUC. We control the fidelity of the training by varying the number of training instances to $\{10\%,25\%,100\%\}$ of the whole training data. For CNN, we additionally set the budget for the number of epochs to $\{10,25,50\}$ and thus obtain a total of $3\times 3$ different levels of fidelity. For FCNet, we vary fidelity by using meta-data at the $\{10,50,100\}$-th epoch. At the end, we obtain a total of $1,500$ landscapes with more than $11$M distinct HP configurations. To further demonstrate the transferability and potential impact of our proposed landscape analysis frameowork, we also employ it to analyze NASBench-101 (Ying et al., 2019), a well-known neural architecture search (NAS) benchmark.
@@ 4 Results and Analysis
In this section, we seek to investigate HP loss landscapes under the four scenarios posed in Section 1. We will start from providing an universal view of the general characteristics of HP loss landscapes of different ML models (Section 4.1). We then explore and compare landscapes under: $i)$ training and test setups (Section 4.2), $ii)$ different fidelities (Section 4.3), $iii)$ different datasets (Section 4.4).
Figure 2: 2D visualization of HP loss landscapes of $6$ ML models under different scenarios: (a) $\mathcal{L}_{\text{test}}$ landscape on baseline datasets (44059 for tree-based models, CIFAR-10 for CNN, protein structure for FCNet), (b) $\mathcal{L}_{\text{train}}$ landscape on baseline datasets, (c) Low-fidelity $\mathcal{L}_{\text{test}}$ landscape on baseline datasets, (d) $\mathcal{L}_{\text{test}}$ landscape on different datasets (44143 for tree-based models, Fashion MINIST for CNN, slice localization for FCNet). Colors indicate $R(\mathcal{L})$ (lower rank values are better).
@@
Figure 3: Distribution of FLA metrics introduced in Table 1 for each model across all datasets for landscapes of 1)1)1 ) $\mathcal{L}_{\text{test}}$, 2)2)2 ) $\mathcal{L}_{\text{train}}$ and 3)3)3 ) $\mathcal{L}_{\text{test}_{\text{LF}}}$.
Figure 4: Distribution of Spearman, Shake-up and $\gamma$-set metrics between (a) $\mathcal{L}_{\text{test}}$ and $\mathcal{L}_{\text{train}}$, (b) $\mathcal{L}_{\text{test}}$ and $\mathcal{L}_{\text{test}_{\text{LF}}}$, (c) $\mathcal{L}_{\text{test}}$ across datasets. Medians are labeled beside each plot.
4.1 Overall Characteristics of HP Loss Landscape of ML Models
From landscape visualizations depicted in Figure 2 (a), we have a general impression that HP loss landscapes of ML models are highly structured and share certain patterns: they are relatively smooth; configurations are clustered in terms of performance; there is a highly distinguishable plateau consisting of prominent configurations, where the terrain becomes flatter. This impression is consistent with the FLA metrics reported in Figure 3, from which we see that landscapes for all models are:
Fairly smooth and clustered. The high $\mathcal{L}$-ast and $\rho_{a}$ values for $\mathcal{L}_{\text{test}}$ landscapes shown in Figure 3 (a) and (b) respectively imply that configurations with similar $\mathcal{L}_{\text{test}}(\bm{\lambda})$ tend to be locally connected, where a small change in $\bm{\lambda}$ is not likely to cause dramatic variation of $\mathcal{L}_{\text{test}}(\bm{\lambda})$. This observation is similar to the findings in reinforcement learning (RL) (Eimer et al., 2023), where the transitions between different parts of the HP landscapes of RL are also found to be quite smooth. This property makes the HP landscapes favorable to Bayesian optimization and search space pruning techniques, as it would be easier to separate the promising regions from the poorly performing ones. On the other hand, if the landscape is rugged instead, in which $\mathcal{L}_{\text{test}}(\bm{\lambda})$ of very different levels often mixed together, it would become more difficult for such techniques to identify a clear promising region.
@@ 4.1 Overall Characteristics of HP Loss Landscape of ML Models
Nearly unimodal. As shown in Figure 3 (e), we find that a considerable fraction of the $\mathcal{L}_{\text{test}}$ landscapes are unimodal, whereas for other landscapes, there could be a handful to dozens (DT) of local optima. This is somewhat contradictory to Pushak & Hoos (2022) at the first thought, in which the authors found that almost all landscapes they studied are nearly unimodal. However, when taking a closer look at the local optima in our landscapes, we find that they usually feature a small basin of attraction (Figure 3 (f)). This makes them relatively ‘shallow’, and thus would not pose significant obstacles to optimization.
However, beyond the results in Figure 3, we find that FCNet landscapes on the $4$ UCI datasets possess $24$ to $347$ local optima, with $\bar{s}_{\mathcal{B}}$ up to $2,372$ (Appendix D), implying a strong attraction for optimizers. Pushak & Hoos have also reported similar observations on these four landscapes, and they speculated the reason could be that these scenarios fall into the over-parameterized regime. While we agree with this reasoning, we seek to conduct further analysis on the local optima using the local optima network (LON) (Ochoa et al. (2008), Appendix B.3). We find that despite the pressence of many other local optima, the global optima still plays a pivotal role in the connectivity of the LON (Appendix D). Therefore, many local optima can eventually escape the global optimum via cetain strategies (e.g., a perturbation), though this may take additional efforts.
@@
Highly neutral; planar around the optimum. As depicted in Figure 3 (d), we can clearly see that HP loss landscapes are often featured in high neutrality. This indicates that a large portion of $1$-bit moves in the landscape will result in subtle change in $\mathcal{L}_{\text{test}}(\bm{\lambda})$ (i.e., $\leq 1\text{\textperthousand}$). We postulate a major reason for this is the low effective dimensionality (Bergstra & Bengio, 2012) of HPO problems: usually only a small subset of all available HPs have obvious influence on performance. Despite landscape neutrality can largely vary with the choice on which HPs to analyze and their respective values, considering the fact that we have ready removed totally unimportant HPs from teh search space, moves with subtle performance shifts can actually be more prevalent than one may expect. Such phenomenon is more pronounced for the well-performing regions, as illustrated by the high NDC values in Figure 3 (c). It suggests that as we move closer to the global optimum, it is more likely to encounter neutral moves, and the landscape becomes flatter. This is in line with Probst & Boulesteix (2017); Pimenta et al. (2020) and practical experience: the gain of tuning HPs usually progressively decreases as approaching the best reachable performance. Such property also poses challenges to optimizers, as there is little gradient information that can be utilized for navigation towards fitter configurations (Muñoz et al., 2015).
Overall, despite certain exceptions, we see that the family of HP loss landscapes tend to share various high-level properties, whereas the exact topologies would vary with models. This explains why in practice, one can often expect an optimizer to work relatively robustly across a wide range of scenarios. In addition, most properties here also seem to generalize the NAS problems (Appendix C), except that we find the NAS landscapes tend to have lower neutrality and more local optima.
4.2 Training and Test HPO Landscapes
@@
Figure 2 (a) and (b) provide a visual comparison of $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes for different models, respectively. We could infer from the plots that the structural characteristics of the $\mathcal{L}_{\text{train}}$ landscapes highly our previously discussed properties. On the other hand, for the performance rankings, we notice that $\mathcal{L}_{\text{train}}$ generally correlate with $\mathcal{L}_{\text{test}}$ for RF, DT, LGBM and CNN, whereas for XGBoost, there is significant shifts in performance between the two cases. We further quantitatively verify such observations using the FLA metrics and the landscape similarity metrics introduced in Section 2.
@@
Structural characteristics. From Figure 3, we could clearly see that the landscape characteristics for $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ are highly consistent for most studied scenarios. More specifically, it is surprising to see that $\mathcal{L}_{\text{train}}$ landscapes tend to yield relatively higher $\mathcal{L}$-ast and $\rho_{a}$, suggesting a smoother and more structured terrain. Meanwhile, the NDC values are lower in the training scenarios. These observations imply that $\mathcal{L}_{\text{train}}$ landscapes are even more benign than $\mathcal{L}_{\text{test}}$ landscapes. In addition, we find that $\bar{\nu}$, $n_{\text{lo}}$ and $\bar{s}_{\mathcal{B}}$ values rarely change between $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes. Notably, local optima found in $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes are almost (if not all) identical. These indicate that the relative performance in local neighborhoods tend to remain similar in the two cases, despite the variations in their numerical values and the global performance rankings.
@@
Landscape similarity in terms of performance rankings. We quantified the similarity between all pairs of $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes for all $5$ models using the three metrics introduced in Section 2 as shown in Figure 4 (a). Overall, we observe that $R(\mathcal{L}_{\text{train}})$ and $R(\mathcal{L}_{\text{test}})$ are globally correlated for all models except XGBoost, as indicated by the significant $\rho_{s}$ values (median $>0.7$) and low Shake-up metrics (median $<0.15$). However, when zooming into the top-$10\%$ regions, we find that the majority of our studied scenarios reveal low $\gamma$-set similarities. It indicates that the generalization gap is larger in prominent regions where configurations are highly adapted to the training set. This phenomenon is more severe for XGBoost, where the median $\gamma$-set similarity is only $0.07$, and there is also a poor $\rho_{s}$ value (median = $0.34$) and high Shake-up score (median = $0.25$).
In order to gain more insight into such generalization gaps for XGBoost, we create scatter plots of $\mathcal{L}_{\text{test}}$ versus $\mathcal{L}_{\text{train}}$ on dataset #44059 as shown in Figure 5 (a). We decompose the pattern into two modes: During the first mode, $\mathcal{L}_{\text{test}}$ highly correlates with $\mathcal{L}_{\text{train}}$ as it decreases, and the models in this stage underfit the data. In the next mode, as points struggle to further move on the $x$-axis ($\mathcal{L}_{\text{train}}$), they stagnate or even significantly increase on the $y$-axis ($\mathcal{L}_{\text{test}}$), indicating strong evidence of overfitting. In particular, we can see a plateauing trend near the $x$-axis, where some models overly excel on the training data, but performing poorly on the test set.
@@
To further investigate which kinds of configurations are likely to lead to overfitting, we color the points with respect to their HP values as shown in Figure 5 (b-e). We are excited to see that the generated plots demonstrate clear patterns between the value of each HP and the resulted performance. In particular, we find that learning rate, max depth and subsample have significant impact on $\Delta\mathcal{L}$. However, the generalizability of a learner is not monopolized by a single one of them; instead, it depends on their cumulative interactions. For example, the largest $\Delta\mathcal{L}$s are observed for learners that features a large learning rate, deep base trees, combined with low subsample rate, but any of these HP settings alone does not necessarily lead to the worst case performance. In addition to this, we notice that such generalization gap is also related to dataset characteristics and weakly correlated across models, and we dissuss more about this matter in Appendix E.
@@
Figure 5: (a) Scatter plot of $\mathcal{L}_{\text{train}}$ versus $\mathcal{L}_{\text{test}}$ for all $14,960$ configurations of XGBoost on the dataset #44059. $\bm{\lambda}^{*}_{\text{test}}$ is marked by red star , (b-e) The same plot with colormap added according to HPs values. Warmer color indicate higher values.
4.3 HPO Landscapes with Different Fidelities

### v2

@@ 3 Experimental Setup
Table 1: Summary of the models and landscape settings included in our empirical study.
Group Model #HPs #Tasks #Fidelities #Configs. #Landscapes
Classical DT $5$ $57$ $\alpha\in\{0.1,0.25,1\}$ $24,200$ $342$
SVM $4$- $\alpha\in\{0.1,0.25,1\}$ $2,340$-
Ensemble XGB $5$ $57$ $\alpha\in\{0.1,0.25,1\}$ $14,960$ $342$
RF $6$ $57$ $\alpha\in\{0.1,0.25,1\}$ $11,250$ $342$
LGBM $5$ $57$ $\alpha\in\{0.1,0.25,1\}$ $13,440$ $342$
DNNs CNN $8$ $6$ $\alpha\in\{0.1,0.25,1\}$, $E\in\{10,25,50\}$ $6,480$ $108$
NB101 NAS $1$ $E\in\{4,12,36,108\}$ $432,624$ $8$
NB201 NAS $3$ $E\in\{10,50,200\}$ $15,625$ $18$
NB360 NAS $2$ $E\in\{10,50,200\}$ $15,625$ $12$
Investigating HPO scenarios posed in Section 1 demands the construction of HP loss landscapes on a wide spectrum of models, tasks and fidelities, and a vast amount of HP configurations. To this end, we conduct large-scale experiments to collect such data, as summarized in Table 1, which consumes over $7$ thousand of GPU and CPU hours (Appendix A) and ends up in $1,500+$ distinct landscapes.
@@
HP search spaces. Due to the presence of continuous HPs, a full enumeration of many HP spaces is not possible. Instead, following [31], we adopt a grid sampling approach to construct HP loss landscapes. Note that while grid search is less efficient than methods like random search in terms of optimization [5], it can yield more interpretable landscapes (e.g., the neighboring configurations only differ in one HP) that are favorable by FLA. Also, though landscape extrapolation via surrogate models can offer infinite sampling resolution [17, 35], they introduce additional biases to the analysis [73]. An important consideration in designing the grid spaces for each model is to ensure that they are representative of the settings commonly used in real-world practice, while keeping a good balance between grid resolution and computational cost. In doing so, we strictly followed our devised protocols in Appendix G.1 to create each space, with typically $2$ to $8$ HPs, and a few to tens of thousands of configurations (Table 1; Appendix G.1). These spaces are not meant to cover the entire feasible region, but rather to focus on more prominent and informative regions as in HPO.
Datasets and configuration evaluation. For the first two groups of models, except SVM, we ground our analysis on $57$ tasks (Appendix G.2) from the tabular benchmark proposed in [67], which span a broad range of complexities (i.e., number of entries and features). For SVM, due to its poor scalability, we are only able to evaluate it on a few tasks. As for CNN, we consider $6$ image classification tasks as detailed in Appendix G.2. For each task, unless predefined, we randomly split the data into training ($80\%$) and test ($20\%$) sets. For all HP configurations $\bm{\lambda}\in\bm{\Lambda}$ of each model, we exhaustively evaluate $\mathcal{L}(\bm{\lambda})_{\text{train}}$ and $\mathcal{L}(\bm{\lambda})_{\text{test}}$ using $5$-fold cross-validation. We use root mean squared error (RMSE) and $R^{2}$ score to serve as the loss function $\mathcal{L}$ for regression tasks, while for classification, we use accuracy and ROC-AUC. We control the fidelity of the training by either varying the ratio of training data ($\alpha$) or the number of training epochs ($E$, Table 1), thus obtaining cheaper proxies of the loss, $\mathcal{L}_{\text{LF}}$.
Distances and neighborhood. After obtaining the set of HP configurations $\bm{\Lambda}$ and their corresponding losses $\mathcal{L}$ on each scenario, we then follow [31] to assess the distance between categorical HPs with the Hamming distance, where each differing value is assigned a $d=1$; for numerical HPs, we use the Manhattan distance, to reflect the number of discrete steps between them on the search grid. The aggregate distance between two configurations, $\bm{\lambda}_{i}$ and $\bm{\lambda}_{j}$, is the sum of these component distances, and we say they are neighbors to each other if $d(\bm{\lambda}_{j},\bm{\lambda}_{i})=1$. Such a neighborhood structure imitates the manual “baby-sitting” tuning process [13] and preserves the local proximity of configurations.
@@ 4 Results and Analysis
We now investigate the obtained HP loss landscapes under the $4$ scenarios posed in Section 1. We begin with exploring the general characteristics of these landscapes in Section 4.1. We then compare landscapes under: $i)$ training and test setups (Section 4.2), $ii)$ different fidelities (Section 4.3), and $iii)$ different datasets (Section 4.4). Note that as we only constructed a few landscapes for SVM, we left its results in Appendix C, where our findings here also generalize to SVM.
Figure 2: $2$D visualization of HP loss landscapes using our proposed method in Section 2 for (A) CNN and (B) XGBoost, w.r.t. $\mathcal{L}_{\mathrm{test}}$ and $\mathcal{L}_{\mathrm{train}}$ as well as different datasets and fidelities (by varying the fraction of training data $\alpha$ or the number of training epochs $E$, marked in red color). (C) shows the visualization of the NASBench-101 landscape with different training epochs on CIFAR-10, and (D) visualizes the NASBench-201 (❶–❸) and NASBench-360 (❹–❺) loss landscapes of different datasets. Colors indicate ranks of performance (lower rank values are better) to normalize the losses.
@@
Figure 3: Distribution of FLA metrics across all datasets for landscapes of 1)1)1 ) $\mathcal{L}_{\mathrm{test}}$, 2)2)2 ) $\mathcal{L}_{\mathrm{train}}$ and 3)3)3 ) $\mathcal{L}_{\mathrm{test}}$ obtained with $10\%$ training data.
4.1 Overall Characteristics of Studied HP Loss Landscapes
We first visualize the topography of the $\mathcal{L}_{\mathrm{test}}$ landscapes for each model using the method in Section 2. Fig. 2 (A-1) and (B-1) show the results for CNN and XGBoost. These provide a general impression of these landscapes, and we clearly observed certain structural similarities: the HP loss landscapes of the studied models are relatively smooth; configurations are locally clustered in terms of their losses; there is a highly distinguishable plateau consisting of prominent configurations, where the terrain becomes flatter. As depicted in Appendix D.1, the landscapes of the other models also exhibit similar patterns. We will then take a closer look at these patterns with quantitative FLA metrics in Fig. 3.
Remark 1. We further verified the reliability of these visualizations in Appendix D.2 with tunable $NK$-landscapes [74] and the black-box optimization benchmark functions from COCO platform [75].
Remark 2. The patterns observed for the systems in Fig. 2 are representative of the other systems, as justified quantitatively by the landscape metrics in Fig. 3 and landscape similarities in Fig. 4.
Fairly smooth and locally clustered. From Fig. 3, we found high loss assortativity and autocorrelation values for the majority of our studied $\mathcal{L}_{\mathrm{test}}$ landscapes. These imply that configurations with similar $\mathcal{L}_{\mathrm{test}}(\bm{\lambda})$ tend to be neighbors to each other, where a small change in $\bm{\lambda}$ is not likely to cause drastic variation in $\mathcal{L}_{\mathrm{test}}(\bm{\lambda})$ (i.e., a fairly smooth landscape). Such property is highly beneficial, as it makes an elegant separation of the prominent regions from the poorly performing ones possible. Such separation can be known as priors, and be utilized to guide the design of search spaces that focus on the promising region [76], or to warm-start HPO solvers directly within that region [53]. Alternatively, this knowledge can also be learned on-the-fly as the optimization proceeds, and thus guides the solver towards fitter areas [4], or assisted in progressively pruning the search space to narrow down the search [77, 78]. On the contrary, if the landscape is highly rugged instead, in which prominent configurations are often interspersed with poorer ones, it would become more difficult to distinguish a clear advantageous region, making it more difficult to navigate the landscape or design a compact search space.
@@ 4.1 Overall Characteristics of Studied HP Loss Landscapes
Nearly unimodal. As shown in Fig. 3 (E), we found a considerable fraction of the $\mathcal{L}_{\mathrm{test}}$ landscapes are unimodal, while for some landscapes (DT, especially), there could be a handful to dozens of local optima. This is somewhat contradictory to [31] at first glance, in which it is found that almost all landscapes studied are nearly unimodal. However, when taking a closer look at the local optima in our landscapes, we found that they are usually featured in a small basin of attraction (see Fig. 3 (F)). This makes them relatively “shallow”, and would not pose significant obstacles to optimization. Thus, we contend that the unimodal hypothesis is still valid for most of our studied scenarios.
@@
Highly neutral; planar around the optimum. As in Fig. 3 (D), we clearly observed that HP loss landscapes are often featured in high neutrality. This indicates that a large portion of purturbations in the landscapes will result in subtle changes in $\mathcal{L}_{\mathrm{test}}$ (i.e., $\leq 1\text{\textperthousand}$). We postulate that a major reason for this is the low effective dimensionality [5] of HPO problems, i.e., usually only a small subset of all tunable HPs have a significant impact on performance. As our designed search spaces have excluded HPs that have no obvious impact on performance, the high neutralities of the resulting landscapes imply that neutral plateaus can actually be more prevalent than one may expect. This phenomenon is more pronounced for the well-performing regions, as illustrated by the high NDC values in Fig. 3 (C). It suggests that as we move closer to the global optimum, it is more likely to encounter neutral moves, and the landscape thus becomes flatter. This observation coincides with experimental observations in HPO, i.e., the gain of tuning HPs often progressively diminishes as approaching higher performance. Such property can pose challenges to the solvers, as there is little gradient information that can be utilized for navigating towards fitter configurations, and they can ‘get lost’ in neutral plateaus. This then underscores the necessity of careful search space design prior to optimization.
Remark 3. The exact measured neutrality of a landscape can depend on multiple factors, including the choice of HPs, the domain, and resolution of the search space, as well as the tolerance $\epsilon$ we used to define neutral moves. Though we have made careful considerations in ensuring the generalizability of our results (see Appendix G), there are no golden standards or consensus for us to explicitly follow. Our goal here is to evoke attention to the potential redundancy in search spaces used in HPO practice, rather than obtaining an absolute measure of neutrality in these landscapes.
The NAS landscapes, as evidenced in Appendix E, also yield high assortativity and autocorrelation values, and the terrain tends to be flatter near the optimum as well. These properties are not surprising as both NAS and HPO are sibling subjects under the umbrella of AutoML. Albeit this, we found that NAS landscapes exhibit considerably more local optima, constituting a multimodal topography (see, e.g., Fig. 2 C-5). In this regard, the NAS problems seem to be more complex. Therefore, we can expect that directly applying HPO solvers to NAS may not lead to satisfactory performance. Instead, tailored search strategies are required to tackle such more rugged landscape surfaces.
Figure 4: Distribution of Spearman, Shake-up and $\gamma$-set metrics between (A) $\mathcal{L}_{\mathrm{test}}$ and $\mathcal{L}_{\mathrm{train}}$, (B) $\mathcal{L}_{\mathrm{test}}$ and $\mathcal{L}_{\mathrm{test}}$ with $10\%$ training data, (C) $\mathcal{L}_{\mathrm{test}}$ across datasets.
4.2 HP Loss Landscapes on Training and Test Data
@@
Fig. 2 (A-2), (B-2) visualize the $\mathcal{L}_{\mathrm{train}}$ landscapes of CNN and XGBoost (see also Appendix D.1). The structural characteristics of these landscapes show high concordance with our previously discussed properties for $\mathcal{L}_{\mathrm{test}}$. On the other hand, for performance rankings, we observed that the $\mathcal{L}_{\mathrm{train}}$ generally correlate with $\mathcal{L}_{\mathrm{test}}$ for CNN, RF, DT, and LGBM, whereas for XGBoost, there is significant discrepancy between the landscapes of $\mathcal{L}_{\mathrm{train}}$ and $\mathcal{L}_{\mathrm{test}}$. We further quantitatively discuss such observations in the following paragraphs.
@@
Landscape characteristics. From Fig. 3, we can clearly see that the structural characteristics for $\mathcal{L}_{\mathrm{train}}$ and $\mathcal{L}_{\mathrm{test}}$ landscapes are highly consistent for most studied scenarios. We found that $\mathcal{L}_{\mathrm{train}}$ landscapes tend to yield relatively higher performance assortativity and autocorrelation values and thus a smoother and more structured terrain. Meanwhile, the NDC values are lower in the training scenarios. These observations imply that the $\mathcal{L}_{\mathrm{train}}$ landscapes are even more benign than $\mathcal{L}_{\mathrm{test}}$ ones. In addition, we found that neutrality, #local optima, and basin sizes rarely change between $\mathcal{L}_{\mathrm{train}}$ and $\mathcal{L}_{\mathrm{test}}$ landscapes. Notably, #local optima found in $\mathcal{L}_{\mathrm{train}}$ and $\mathcal{L}_{\mathrm{test}}$ landscapes are almost (if not all) identical. This suggests that the relative performance in local neighborhoods tends to be retained in the two cases, despite variations in their numerical values and global rankings.
@@
Landscape similarity. We quantified the similarity between all pairs of $\mathcal{L}_{\mathrm{train}}$ and $\mathcal{L}_{\mathrm{test}}$ landscapes using the $3$ metrics introduced in Section 2 as shown in Fig. 4 (A). Overall, we observed that $R(\mathcal{L}_{\mathrm{train}})$ and $R(\mathcal{L}_{\mathrm{test}})$ are globally correlated for all models except XGBoost, as indicated by the significant $\rho_{s}$ values (median $>0.7$) and low Shake-up metrics (median $<0.15$). However, when zooming into the top-$10\%$ regions, we saw that the majority of our studied scenarios reveal low $\gamma$-set similarities. It indicates that the generalization gap is larger in prominent regions where configurations are highly adapted to the training set. This is more severe for XGBoost, where the median $\gamma$-set similarity is only $0.07$, and there is also a low $\rho_{s}$ value (median = $0.34$) and high Shake-up score (median = $0.25$).
@@
A closer look at overfitting. In order to gain more insight into such generalization gaps for XGBoost, we create scatter plots of $\mathcal{L}_{\mathrm{test}}$ versus $\mathcal{L}_{\mathrm{train}}$ on the dataset #44059 in Fig. 5 (A), where we can decompose the pattern into two modes. In the first one, $\mathcal{L}_{\mathrm{test}}$ highly correlates with $\mathcal{L}_{\mathrm{train}}$ as it decreases, and the models in this stage underfit the data. In the second mode, on the contrary, as points struggle to further move on the $x$-axis ($\mathcal{L}_{\mathrm{train}}$), they stagnate or even significantly increase on the $y$-axis ($\mathcal{L}_{\mathrm{test}}$), indicating strong evidence of overfitting. In particular, we can see a plateauing trend near the $x$-axis, where some models overly excel on the training data, but perform poorly on the test set. To further investigate which kinds of configurations are likely to lead to overfitting, we color the points with respect to their HP values as shown in Fig. 5 (B-E). It is interesting to see that the generated plots demonstrate clear patterns between the value of each HP and the resulting performance. In particular, we find that learning rate, max depth, and subsample have a significant impact on $\Delta\mathcal{L}$. However, the generalizability of a learner is not monopolized by a single one of them; instead, it depends on their cumulative interactions. For example, the largest $\Delta\mathcal{L}$s are observed for learners that feature a large learning rate, and deep base trees, combined with a low subsample rate, but any of these HP settings alone does not necessarily lead to the worst case performance. In addition to this, we noticed that such a generalization gap is also correlated some dataset properties such as number of samples or features, and we discuss more about this in Appendix F.
@@
Figure 5: (A) Scatter plot of $\mathcal{L}_{\mathrm{train}}$ versus $\mathcal{L}_{\mathrm{test}}$ of XGBoost on the dataset #44059. $\bm{\lambda}^{*}_{\mathrm{test}}$ is marked by , (B-E) The same plot with colors to indicate HPs values. Warmer color indicate higher values.
4.3 HP Loss Landscapes with Different Fidelities